OpenStudy (anonymous):

2[7-5(1-n)]+5n=-6+2[6(n+1)-3n]

OpenStudy (yanasidlinskiy):

Use P.E.M.D.A.S.

OpenStudy (anonymous):

Hello, this is not the exact question, but this shows you the exact steps on how to complete this problem. :D Simplifying 2[7 + -5(1 + -1n)] + 5n = -18 + 4[6(n + 1) + -3n] 2[7 + (1 * -5 + -1n * -5)] + 5n = -18 + 4[6(n + 1) + -3n] 2[7 + (-5 + 5n)] + 5n = -18 + 4[6(n + 1) + -3n] Combine like termsAd by save keep: 7 + -5 = 2 2[2 + 5n] + 5n = -18 + 4[6(n + 1) + -3n] [2 * 2 + 5n * 2] + 5n = -18 + 4[6(n + 1) + -3n] [4 + 10n] + 5n = -18 + 4[6(n + 1) + -3n] Combine like terms: 10n + 5n = 15n 4 + 15n = -18 + 4[6(n + 1) + -3n] Reorder the terms: 4 + 15n = -18 + 4[6(1 + n) + -3n] 4 + 15n = -18 + 4[(1 * 6 + n * 6) + -3n] 4 + 15n = -18 + 4[(6 + 6n) + -3n] Combine like terms: 6n + -3n = 3n 4 + 15n = -18 + 4[6 + 3n] 4 + 15n = -18 + [6 * 4 + 3n * 4] 4 + 15n = -18 + [24 + 12n] Combine like terms: -18 + 24 = 6 4 + 15n = 6 + 12n Solving 4 + 15n = 6 + 12n Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '-12n' to each side of the equation. 4 + 15n + -12n = 6 + 12n + -12n Combine like terms: 15n + -12n = 3n 4 + 3n = 6 + 12n + -12n Combine like terms: 12n + -12n = 0 4 + 3n = 6 + 0 4 + 3n = 6 Add '-4' to each side of the equation. 4 + -4 + 3n = 6 + -4 Combine like terms: 4 + -4 = 0 0 + 3n = 6 + -4 3n = 6 + -4 Combine like terms: 6 + -4 = 2 3n = 2 Divide each side by '3'. n = 0.6666666667 Simplifying n = 0.6666666667 @jessicaaaaaray!

OpenStudy (anonymous):

I've used PEMDAS, but keep getting the wrong answer

OpenStudy (yanasidlinskiy):

Umm...That isn't the right answer...(I do know that this question is different.) @jessicaaaaaray! Give me some time to solve it on my end..Then I'll show you how...If this is a multiple choice question with choices abcd....Share them with me here..

OpenStudy (anonymous):

I'm sorry it's not multiple choice. Thank you though!

OpenStudy (yanasidlinskiy):

Ok. Before I start typing this out..I'll let you know this might take a while...sorry..:(

OpenStudy (anonymous):

Thank you though!

OpenStudy (yanasidlinskiy):

2[7-5(1-n]+5n=-6+2[6n+1-3n] 2[7-5+5n]+5n=6+2[3n+1 2[2+5n]+5n=6+6n+2 4+10n+5n=-6+6+3 --------Combine variables on both sides! 4+15n=-4+6n -------Now let me ask you to solve it from there. Can you?

OpenStudy (anonymous):

Yeah, I got it. 2/9 ?